# Informationally overcomplete measurements from generalized equiangular tight frames

### v2

@@ 2 Generalized equiangular measurements
$\frac{1}{d}<b_{\alpha}\leq\frac{1}{d}\max\{d,M_{\alpha}\}.$ (11)
@@ 4 Conical designs
$0<S\leq\min\left\{\frac{\gamma_{\alpha}^{2}}{M_{\alpha}},\frac{d-1}{M_{\alpha}-1}\frac{\gamma_{\alpha}^{2}}{M_{\alpha}}\right\}$ (30)
@@ 5 Index of coincidence
Now, $\{p_{\alpha,k}=\mathrm{Tr}(\rho P_{\alpha,k});\,k=1,\ldots,M_{\alpha};\,\alpha=1,\ldots,N\}$ is a probability distribution, whereas $Q_{\alpha,k}$ is a dual frame to $P_{\alpha,k}$ defined by $\mathrm{Tr}(P_{\alpha,k}Q_{\beta,\ell})=\delta_{\alpha\beta}\delta_{k\ell}$. Due to overcompleteness of the generalized equiangular measurements, the choice of a dual frame is not unique. However, for our purposes, we distinguish a special frame: the one that can be constructed from
@@ Appendix B Conical designs
$_{\alpha}\quad S\leq\frac{\gamma_{\alpha}^{2}}{M_{\alpha}}\qquad{\rm and}\qquad\mbox{\Large$\forall$}_{\alpha}\quad S\leq\frac{d-1}{M_{\alpha}-1}\frac{\gamma_{\alpha}^{2}}{M_{\alpha}}.$ (52)

### v3

@@ 2 Generalized equiangular measurements
$\frac{1}{d}<b_{\alpha}\leq\frac{1}{d}\min\{d,M_{\alpha}\}.$ (11)
@@ 4 Conical designs
$0<S\leq\min\left\{\frac{d\gamma_{\alpha}^{2}}{M_{\alpha}},\frac{d-1}{M_{\alpha}-1}\frac{d\gamma_{\alpha}^{2}}{M_{\alpha}}\right\}$ (30)
@@ 5 Index of coincidence
Now, $\{p_{\alpha,k}=\mathrm{Tr}(\rho P_{\alpha,k});\,k=1,\ldots,M_{\alpha};\,\alpha=1,\ldots,N\}$ is a probability distribution, whereas $Q_{\alpha,k}$ is a dual frame to $P_{\alpha,k}$. Due to overcompleteness of the generalized equiangular measurements, the choice of a dual frame is not unique. However, for our purposes, we distinguish a special frame: the one that can be constructed from
@@ Appendix B Conical designs
$_{\alpha}\quad S\leq\frac{d\gamma_{\alpha}^{2}}{M_{\alpha}}\qquad{\rm and}\qquad\mbox{\Large$\forall$}_{\alpha}\quad S\leq\frac{d-1}{M_{\alpha}-1}\frac{d\gamma_{\alpha}^{2}}{M_{\alpha}}.$ (52)